# Josephson traveling-wave parametric amplifier based on low-intrinsic-loss coplanar lumped-element waveguide

### v2

@@ Appendix G FABRICATION RECIPE
G.1 Part 1: Base circuit fabrication by ultraviolet lithography
G.1.1 Substrate preparation
@@ G.1.3 Pattern transfer
G.1.4 Resist removal and final cleaning
@@
G.2 Part 2: Junction fabrication by electron-beam lithography
G.2.1 Bilayer resist application
@@
G.2.2 Electron-beam exposure and development
@@
G.2.3 Junction evaporation and liftoff
Josephson junctions are formed by double-angle shadow evaporation, with $>$3-hour pump-down to below 10-7 Torr. We avoid argon-ion milling to prevent contamination of the silicon substrate by resist sputtering [dunsworthCharacterizationReductionCapacitive2017a]. The first layer deposition consists of 2-minute titanium gettering at 0.1 nm/s, 30-nm aluminum electrode at 0.2 nm/s, and 8-minute oxidation at 0.12 Torr (yielding $\approx$75 $\Omega$/µm2). The second layer includes 2-minute titanium gettering, 90-nm aluminum at 0.2 nm/s, and 5-minute oxidation at 15 Torr for passivation.
@@ G.2.3 Junction evaporation and liftoff
G.3 Part 3: Airbridge fabrication and final processing
G.3.1 Support layer formation
@@ G.3.2 Airbridge metallization
G.3.3 Pattern definition and etching
@@
G.3.4 Final dicing and cleaning

### v3

@@ Appendix H FABRICATION RECIPE
H.1 Part 1: Base circuit fabrication by ultraviolet lithography
H.1.1 Substrate preparation
@@ H.1.3 Pattern transfer
H.1.4 Resist removal and final cleaning
@@
H.2 Part 2: Junction fabrication by electron-beam lithography
H.2.1 Bilayer resist application
@@
H.2.2 Electron-beam exposure and development
@@
H.2.3 Junction evaporation and liftoff
Josephson junctions are formed by double-angle shadow evaporation, with >>3-hour pump-down to below 10-7 Torr. We avoid argon-ion milling to prevent contamination of the silicon substrate by resist sputtering [50]. The first layer deposition consists of 2-minute titanium gettering at 0.1 nm/s, 30-nm aluminum electrode at 0.2 nm/s, and 8-minute oxidation at 0.12 Torr (yielding $\approx$75 $\Omega$/µm2). The second layer includes 2-minute titanium gettering, 90-nm aluminum at 0.2 nm/s, and 5-minute oxidation at 15 Torr for passivation.
@@ H.2.3 Junction evaporation and liftoff
H.3 Part 3: Airbridge fabrication and final processing
H.3.1 Support layer formation
@@ H.3.2 Airbridge metallization
H.3.3 Pattern definition and etching
@@
H.3.4 Final dicing and cleaning